Solve the following system of linear equations by graphing. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
18x-3y=3
6x+y=-1
x = ?
y = ?
pls show me the steps to get the answer( not by graph)

Answer by mananth(16946)

18x-3y=3 .............1
6x+y=-1 .............2

multiply (1)by 1
Multiply (2) by 3
18x-3y= 3
18x+3y= -3
Add the two equations
36x=0
x=0
plug value of x in (1)

18x-3y=3
-3y=3
3y=3
-3y=3
y=-1